7. Constant growth rates One of the most important components of stock valuation is a firms estimated growth rate. FinancialBased on this information, Roberts forecast of PAMCs growth rate in earnings and dividends should be: 08.15% o 7.35% 0 28.7
Answer #1

Return on Equity=Net Income/Book Value of Equity=1341300/10497500=12.77733%

Growth rate=Return on Equity*(1-Dividend Payout Rate)=12.77733%*(1-42.5%)=7.34696%

Retaining a higher percentage of earnings will result in a higher growth rate